OpenStudy (anonymous):

If f(x) = x + 2, find f(3). 1 6 5 −1

OpenStudy (zzr0ck3r):

f(3) = 3+2 you got it from here?

OpenStudy (unklerhaukus):

\[f(x)=(x)+2\] \[f(3)=(3)+2\]\[\qquad=3+2=\dots\]

OpenStudy (anonymous):

5???

OpenStudy (zzr0ck3r):

yes